# Effective Hamiltonian approach to the Quantum Phase transitions in the extended Jaynes-Cummings model

### 11 Dec 2023 v2

@@ Abstract
The study of phase transitions in dissipative quantum systems based on the Liouvillian is often hindered by the difficulty of constructing a time-local master equation when the system-environment coupling is strong. To address this issue, the complex discretization approximation for the environment has been proposed to study the quantum phase transition in the extended Jaynes-Cumming model with an infinite number of boson modes. This approach yields a non-Hermitian effective Hamiltonian that can be used to simulate the dynamics of the spin. It was found that the ground state of this effective Hamiltonian determines the spin dynamics in the single-excitation subspace. Depending on the opening of the energy gap and the maximum population of excitations on the spin degree of freedom, three distinct phases can be identified: fast decaying, localized, and stretched dynamics of the spin. This approach can be extended to multiple excitations, and similar dynamics were found in the double-excitation subspace, indicating the robustness of the single-excitation phase.
@@
The study of open quantum systems has revealed that phase transitions can occur, even though the system has a finite number of degrees of freedom hwang2016 . Unlike the phase transitions in closed systems under thermodynamic conditions, the phase transitions in open quantum systems are accompanied by significant changes in the system’s dissipative dynamics. Thus, the so called dissipative phase transition (DPT) reflects the non-equilibrium of system and the nontrivial properties of steady state of the system. Moreover by environment engineering, one can prepare the system in a special state through DPT, that would be difficult or impossible by conventional ways diehl2008 . To characterize DPT explicitly, the system’s Liouvillian eigenvalues is studied , of which the real part corresponds to the decay rate of eigenmodes kessler12 . It was noted that the vanishing of the gap in the excitation spectrum of Liouvillian indicates a non-analytical change in the dissipative dynamics of system, which typically signals the onset of a phase transition kessler12 .
@@
Recent experimental advances have allowed physicists to investigate regimes beyond Markovian approximation, such as in solid-state solid-state and artificial light-matter systems light-matter . In this situation, the Liouvillian is difficult to determine due to the memory effect of environment. This raise the question of whether phase transitions can still be found in the absence of a Liouvillian and if so, how to characterize it explicitly. To address this, the method of perturbational expansion or numerical simulation has been used to incorporate the influence of environment vega2017 . However, the resulting dynamical equations are often very complicated, making it difficult to determine the phase transition breuer2002 . An alternative approach is to approximate the environment as a many-body system with discrete and finite energy modes. Thus, an effective Hamiltonian can be constructed to predicate the dynamics of system kazansky97 ; discretization . However, this approach requires a large Hilbert space dimension to accurately predicate the long-term behavior of system, which would become unaffordable in computation. To surmount this difficulty, a complex discretization approximation for environment was proposed by the authors cui23 , which allows for a more efficient computation and more faithful depiction of the dissipative dynamics of system by introducing an effective Hamiltonian. Thanking to the non-Hermitianity of effective Hamiltonian, the decay rate of eigenmode can be characterized by the imaginary part of corresponding eigenvalue.
@@
The similarity of the Liouvillian and the effective Hamiltonian encourages us to explore the phase transition by solving the complex Hamiltonian. This approach incorporates the dynamical information for both the system and environment. So by studying the eigenvalues and corresponding eigenvectors, one can gain the deep insight into the dynamics of system. As an illustration, the extended Jaynes-Cumming model is discussed in this paper. This model is the generalization of the spin-boson model under rotating-wave approximation. It was known that the spin-boson model undergoes a quantum phase transition in the sub-Ohmic or Ohmic regime from a delocalized phase with a single ground state and no magnetization, to a localized phase with two-fold degenerate ground state and a nonzero magnetization leggett . However, the nonpreservation of total excitations in the spin-boson model makes the investigation of the phase transition complicated sbmodel . The rotating-wave approximation is a reasonable choice to obtain tractable solutions breuer2002 . Furthermore, it allows us to determine the phase transition by finding singularity of order parameter. A similar research has been performed in Ref. wang19 , where the effective Hamiltonian is constructed in real space. Thus, the magnetization $\langle\sigma_{z}\rangle$ and $\langle\sigma_{x}\rangle$ are evaluated for the ground state to detect phase transitions.
@@
In constrast, a non-Hermtian effective Hamiltonian is constructed in this paper to determine the phase transition in the single-excitation subspace of extended Jaynes-Cumming model. Because of the non-Hermitianity, the eigenvalue of effective Hamiltonian shows the imaginary part, which characterizes the decaying dynamics of spin. Importantly, besides of the known dacaying phase and localization phase kofman1994 , the intermediate phase can be verified through checking the ground state function. The intermediate phase demonstrates stretched-like behavior in dynamics. By our knowledge, it is the first time for the identification of this phase in the extended Jaynes-Cumming model. Moreover, the approach can be readily extended into the multi-excitation subspace. As a result, we explore the spin dynamics in the double-excitation subspace by exact digitalization method. Interestingly, the spin dynamics in this case could shows the similar behavior as that in single-excitation subspace, which implies that the single-excitation phases would be stable.
@@
The remaining of the paper is organized as follows. The model and method are introduced in Section II, with particular remarks on the complex discretization approximation for environment. In section III, the spin dynamics is discussed explicitly in the single-excitation subspace by solving the effective complex Hamiltonian. Resultantly, a phase diagram is founded. In section IV, we extend the discussion into the double-excitation subspace. It is found that the spin dynamics in this case is nearly the same as in single-excitation subspace. Finally, conclusions are offered in section V.
@@ II Model and method
In this section, we introduce the model and present the overview of complex discretization approximation.
@@ II.1 Extended Jaynes-Cumming model with infinite number of boson modes
where $\sigma_{\pm}=\left(\sigma_{x}\pm\mathbbm{i}\sigma_{y}\right)/2$ with $\sigma_{i}(i=x,y,z)$ the Pauli matrices and $\mathbbm{i}$ is the imaginary unit. $a^{\dagger}_{k}(a_{k})$ denotes the creation (annihilation) operator of mode $k$ in environment, which is coupled to the spin via coupling strength $g_{k}$. In this paper, the spectral density for environment is chosen as
$J(\omega)=\sum_{k}\left|g_{k}\right|^{2}\delta\left(\omega-\omega_{k}\right)=\eta\omega\left(\frac{\omega}{\omega_{c}}\right)^{s-1}e^{-x/\omega_{c}},$ (2)
where $\eta$ depicts the coupling strength, and $\omega_{c}$ is the cut-off of frequency. Depending the value of $s$, the environment is classified as sub-Ohmic $\left(0<s<1\right)$, Ohmic $\left(s=1\right)$ and super-Ohmic $\left(s>1\right)$. Although the simplicity of Eq. (1), the exact spin dynamics can be obtained only in the single-excitation subspace. In this case, the state of total system can be written as
@@
in which $N$ denotes the number of modes in environment, $|1\rangle_{k}=a^{\dagger}_{k}|0\rangle_{k}$ with the vacuum state $|0\rangle_{k}$, and $|e\rangle=\sigma_{+}|g\rangle$. Substituting Eq. (3) into the Schrödinger equation, one gets
@@
where the last term stems from the memory effect of environment.
It was known that the spin dynamics in the single-excitation subspace can display the transition from decaying to localization because of the occurrence of bound state kofman1994 ; bs . The bound state corresponds to a singular energy level, separated from the continuum of environment. The existence of bound staten can be verified by solving Eq. (4) kofman1994 . By Laplace transformation, Eq. (4) can be rewritten as
@@
$\Delta-E-\int_{0}^{\infty}\text{d}\omega\frac{J(\omega)}{E-\omega}=0,$ (6)
one then obtains $\alpha(t)\propto e^{-\mathbbm{i}Et}$ by inverse Laplace transformation. This feature means that the modulus of $\alpha(t)$ may be constant, leading to localization in the ynamics. The real solution to Eq. (6) can be found in the regime of $E<0$, depending on $J(\omega)$ bs . Moreover, since $\omega_{k}\in\left[0,\infty\right)$, the bound state can be viewed as the ground state of total system. In this sense, the transition is a result of the opening of energy gap of ground state from the continuum.
Eq. (4) can be solved using iterative method, but the memory effect renders the calculation time-consuming exhaustive for long term evolution. For multiple excitations, the lack of a dynamical equation similar to Eq. (4) renders it difficult to analyze the spin dynamics and determine the bound state shi16 . To address this, a feasible approach is to discretize the environment into a finite set of modes,thereby constructing an effective Hamiltonian that can be solved with exact diagonalization discretization . However, this approach suffers from the recurrence caused by the finite dimension, which make the evaluation unaffordable for long term evolution.
@@ II.2 Complex discretization approximation
Figure 2: (Color online)The plots for the computational errors defined in Eq. (21). For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
Figure 3: (Color online) The plot for the computational errors when $\widetilde{H}_{\text{dis}}$ is used to evaluate the survival probability for $s=1$. It is stressed that the data for $\eta=0.05$ and $0.1$ has been enlarged by ten times in convenience of illustration. The other parameters are the same as those in Fig. 2.
Recently, the complex discretization approximation (CDA) method has been introduced to refine the simulation of spin dynamics cui23 . In contrast to its counterpart in real space discretization , this method yields a non-Hermitian effective Hamiltonian $H_{\text{dis}}$, which has complex eigenvalues with a negative imaginary part. Consequently, $H_{\text{dis}}$ can provide more accurate description of the dissipative dynamics of spin.
The discretization approximation was proposed to tackle the infinite energy modes in environment by utilizing the Gauss quadrature discretization . For this purpose, Eq. (1) is transformed under the continuous limit as bulla1997 ,
@@
$\int_{0}^{\infty}\text{d}x\left[g(x)a^{\dagger}_{x}\sigma_{-}+g*(x)a_{x}\sigma_{+}\right].$
@@
The equivalence between Eq. (1) and Eq. (7) is guaranteed by the fact that the same effective action for spin degree of freedom can be derived from the Hamiltonians Eq. (1) and Eq. (7) bulla1997 .
It is noteworthy that the forms of $h(x)$ or $g(x)$ are not uniquely determined for a given $J(x)$. By this freedom, one can choose proper form for $h(x)$ or $g(x)$ for the sake of discretization of environment discretization . Given Ohmic spectral function Eq. (2), one may choose
@@
The main idea of CDA is to utilize complex Gauss quadratures to discretize continuum of the environment. For this purpose, $\int_{0}^{\infty}\text{d}x$ is replaced by a line integral $\int_{\Gamma}\text{d}z$ in complex plane shown in Fig. 1, where $z=x+\mathbbm{i}y$. $\Gamma$ denotes the semicircle with radius $R$ in the lower half complex plane, centered at $\left(R,0\right)$. With the requirement that both $h(x)$ and $g(x)$ have no singularity in complex plane, $\int_{\Gamma}\text{d}z$ is equal to $\int_{0}^{\infty}\text{d}x$ by Cauchy theorem if $R\rightarrow\infty$. Along the path $\Gamma$, $z=R(1+e^{\mathbbm{i}\theta})$ with $\theta\in\left[\pi,2\pi\right]$. Thus, the real part of $z$ is ensured to be nonnegative, characterizing the energy mode in environment. The value of $R$ is responsible for the effective energy modes in environment, used to simulate the spin dynamics.
Similar to the way in discretization , the following transformation is introduced
@@
It is stressed that the inner product is defined deliberately without complex conjugation to construct the three-term recurrence relation for $\eta_{n}(z)$ gautschi . By this way, $\eta_{n}(z)$ can display two features, which are crucial for discretization of Eq.7. One is the orthonormality defined as
@@
where $N$ is the highest degree of polynomial, used in the evaluation. $h(z)$ and $g(z)$ can obtained directly through replacing $x$ in Eq. (9) by $z$.
As further simplification, let introduce the new mode operator
@@
is presented in Fig. 2 for different values of $s$ and $\eta$ (the explicit plots for the evolution of survival probability are provided by Fig.A1 in Appendix I). Nonetheless, due to the time-consuming nature of solving Eq. (4), the exact evaluation are restricted up to $t=10$ with the step length of $10^{-4}$. For CDA, the weight function chosen is $w(z)=1$, which might cause the appearance of complex eigenmodes in the environment. It should be stressed that choosing $g(z)$ as the weight function will not generate a non-Hermitian effective Hamiltonian. Evidently, the computational errors has a magnitude of $10^{-3}$ in most circumstances, and thus CDA can provides reliable evaluation of spin dynamics.
However, Fig. 2(c) shows a significant increase of error when $s=1$. The explicit calculations reveals a slow convergence to stable result as $N$ increases. To capture the long-term behavior of spin dynamics, we propose to use

### 06 Apr 2024 v3

@@ Abstract
The study of phase transitions in dissipative quantum systems based on the Liouvillian is often hindered by the difficulty of constructing a time-local master equation when the system-environment coupling is strong. To address this issue, the complex discretization approximation for the environment is proposed to study the quantum phase transition in the extended Jaynes-Cumming model with an infinite number of boson modes. This approach yields a non-Hermitian effective Hamiltonian that can be used to simulate the dynamics of the spin. It is found that the ground state of this effective Hamiltonian determines the spin dynamics in the single-excitation subspace. Depending on the opening of the energy gap and the maximum population of excitations on the spin degree of freedom, three distinct phases can be identified: fast decaying, localized, and stretched dynamics of the spin. This approach can be extended to multiple excitations, and similar dynamics were found in the double-excitation subspace, indicating the robustness of the single-excitation phase.
@@
The study of open quantum systems has revealed that phase transitions can occur, even though the system has a finite number of degrees of freedom hwang2016 . Unlike the phase transitions in closed systems under thermodynamic conditions, the phase transitions in open quantum systems are accompanied by significant changes in the system’s dissipative dynamics. Thus, the so called dissipative phase transition (DPT) reflects the non-equilibrium of a system and the nontrivial properties of the steady state of the system. Moreover, by environment engineering, one can prepare the system in a special state through DPT, that would be difficult or impossible by conventional ways diehl2008 . To characterize DPT explicitly, the system’s Liouvillian eigenvalues are studied , of which the real part corresponds to the decay rate of eigenmodes kessler12 . It was noted that the vanishing of the gap in the excitation spectrum of Liouvillian indicates a non-analytical change in the dissipative dynamics of system, which typically signals the onset of a phase transition kessler12 .
@@
Recent experimental advances have allowed physicists to investigate regimes beyond Markovian approximation, such as in solid-state solid-state and artificial light-matter systems light-matter . In this situation, the Liouvillian is difficult to determine due to the memory effect of the environment. This raise the question of whether phase transitions can still be found in the absence of a Liouvillian and if so, how to characterize it explicitly. To address this, the method of perturbational expansion or numerical simulation has been used to incorporate the influence of the environment vega2017 . However, the resulting dynamical equations are often very complicated, making it difficult to determine the phase transition breuer2002 . An alternative approach is to approximate the environment as a many-body system with discrete and finite energy modes. Thus, an effective Hamiltonian can be constructed to predicate the dynamics of system kazansky97 ; discretization . However, this approach requires a large Hilbert space dimension to accurately predicate the long-term behavior of system, which would become unaffordable in computation. To surmount this difficulty, a complex discretization approximation for the environment was proposed by the authors cui23 , which allows for a more efficient computation and more faithful depiction of the dissipative dynamics of system by introducing an effective Hamiltonian. Thanking to the non-Hermitianity of effective Hamiltonian, the decay rate of the eigenmode can be characterized by the imaginary part of the corresponding eigenvalue.
@@
The similarity of the Liouvillian and the effective Hamiltonian encourages us to explore the phase transition by solving the complex Hamiltonian. This approach incorporates the dynamical information for both the system and environment. So by studying the eigenvalues and corresponding eigenvectors, one can gain deep insight into the dynamics of a system. As an illustration, the extended Jaynes-Cumming model is discussed in this paper. This model is the generalization of the spin-boson model under rotating-wave approximation. It was known that the spin-boson model undergoes a quantum phase transition in the sub-Ohmic or Ohmic regime from a delocalized phase with a single ground state and no magnetization, to a localized phase with two-fold degenerate ground state and a nonzero magnetization leggett . However, the nonpreservation of total excitations in the spin-boson model makes the investigation of the phase transition complicated sbmodel . The rotating-wave approximation is a reasonable choice to obtain tractable solutions breuer2002 . Furthermore, it allows us to determine the phase transition by finding the singularity of order parameter. Similar research was performed in Ref. wang19 , where the effective Hamiltonian was constructed in real space. Thus, the magnetization $\langle\sigma_{z}\rangle$ and $\langle\sigma_{x}\rangle$ are evaluated for the ground state to detect phase transitions.
@@
In contrast, a non-Hermtian effective Hamiltonian is constructed in this paper to determine the phase transition in the single-excitation subspace of the extended Jaynes-Cumming model. Because of the non-Hermitianity, the eigenvalue of the effective Hamiltonian shows the imaginary part, which characterizes the decaying dynamics of spin. Importantly, besides the known decaying phase and localization phase kofman1994 , the intermediate phase can be verified through checking the ground state function. The intermediate phase demonstrates stretched-like behavior in dynamics. The approach can be readily extended into the multi-excitation subspace. As a result, we explore the spin dynamics in the double-excitation subspace by the exact diagonalizatin method. Interestingly, the spin dynamics in this case could show similar behavior as that in the single-excitation subspace, which implies that the single-excitation phases would be stable.
@@
The remainder of the paper is organized as follows. The model and method are introduced in Sec. II, with particular remarks on the complex discretization approximation for environment. In Sec. III, the spin dynamics is discussed explicitly in the single-excitation subspace by solving the effective complex Hamiltonian. Resultantly, a phase diagram is founded. In Sec. IV, we extend the discussion into the double-excitation subspace. It is found that the spin dynamics in this case is nearly the same as in single-excitation subspace. Finally, conclusions are offered in Sec. V.
@@ II Model and method
In this section, we introduce the model and present the overview of the complex discretization approximation.
@@ II.1 Extended Jaynes-Cumming model with infinite number of boson modes
where $\sigma_{\pm}=\left(\sigma_{x}\pm\mathbbm{i}\sigma_{y}\right)/2$ with $\sigma_{i}(i=x,y,z)$ the Pauli matrices and $\mathbbm{i}$ is the imaginary unit. $a^{\dagger}_{k}(a_{k})$ denotes the creation (annihilation) operator of mode $k$ in the environment, which is coupled to the spin via coupling strength $g_{k}$. In this paper, the spectral density for the environment is chosen as
$J(\omega)=\sum_{k}\left|g_{k}\right|^{2}\delta\left(\omega-\omega_{k}\right)=\eta\omega\left(\frac{\omega}{\omega_{c}}\right)^{s-1}e^{-\omega/\omega_{c}},$ (2)
where $\eta$ depicts the coupling strength, and $\omega_{c}$ is the cut-off of frequency. Depending on the value of $s$, the environment is classified as sub-Ohmic $\left(0<s<1\right)$, Ohmic $\left(s=1\right)$ or super-Ohmic $\left(s>1\right)$. Despite the simplicity of Eq. (1), the exact spin dynamics can be obtained only in the single-excitation subspace. In this case, the state of the total system can be written as
@@
in which $N$ denotes the number of modes in the environment, $|1\rangle_{k}=a^{\dagger}_{k}|0\rangle_{k}$ with the vacuum state $|0\rangle_{k}$, and $|e\rangle=\sigma_{+}|g\rangle$. Substituting Eq. (3) into the Schrödinger equation, one gets
@@
where the last term stems from the memory effect of the environment.
It is known that the spin dynamics in the single-excitation subspace can display the transition from decaying to localization because of the occurrence of a bound state kofman1994 ; bs . The bound state corresponds to a singular energy level, separated from the continuum of the environment. The existence of bound staten can be verified by solving Eq. (4) kofman1994 . By using Laplace transformation, Eq. (4) can be rewritten as
@@
$\Delta-E+\int_{0}^{\infty}\text{d}\omega\frac{J(\omega)}{E-\omega}=0,$ (6)
one then obtains $\alpha(t)\propto e^{-\mathbbm{i}Et}$ by inverse Laplace transformation. This feature means that the modulus of $\alpha(t)$ may be constant, leading to localization in the dynamics. The real solution to Eq. (6) can be found in the regime of $E<0$, depending on $J(\omega)$ bs . Moreover, since $\omega_{k}\in\left[0,\infty\right)$, the bound state can be viewed as the ground state of the total system. In this sense, the transition is a result of the opening of energy gap of the ground state from the continuum.
Equation (4) can be solved by using the iterative method, but the memory effect renders the calculation time-consuming and exhaustive for long term evolution. For multiple excitations, the lack of a dynamical equation similar to Eq. (4) renders it difficult to analyze the spin dynamics and determine the bound state shi16 . To address this, a feasible approach is to discretize the environment into a finite set of modes, thereby constructing an effective Hamiltonian that can be solved with exact diagonalization discretization . However, this approach suffers from the recurrence caused by the finite dimension, which make the evaluation unaffordable for long term evolution.
@@ II.2 Complex discretization approximation
Figure 2: (Color online)The plots for the computational errors defined in Eq. (21) vs. the evolution time $t$ (in units of $1/\Delta$). For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
Figure 3: (Color online)(a) The plot for the computational errors vs. the evolution time $t$ (in units of $1/\Delta$) when $\widetilde{H}_{\text{dis}}$ is used to evaluate the survival probability for $s=1$. It is stressed that the data for $\eta=0.05$ and $0.1$ are enlarged by ten times for convenience of illustration. The other parameters are the same as those in Fig. 2. (b) The detailed plot for the oscillation of survival probability when $\eta=0.5$. The dashed line denotes the value $0.54744$. The evaluation is obtained by $\widetilde{H}_{\text{dis}}$ with $N=2000$. The other parameters are the same as those in Fig. 2.
Recently, the complex discretization approximation (CDA) method was introduced to refine the simulation of spin dynamics cui23 . In contrast to its counterpart in real space discretization , this method yields a non-Hermitian effective Hamiltonian $H_{\text{dis}}$, which has complex eigenvalues with a negative imaginary part. Consequently, $H_{\text{dis}}$ can provide a more accurate description of the dissipative dynamics of spin.
The discretization approximation was proposed to tackle the infinite energy modes in an environment by utilizing the Gauss quadrature discretization . For this purpose, Eq. (1) is transformed under the continuous limit as bulla1997 ,
@@
$\int_{0}^{\infty}\text{d}x\left[g(x)a^{\dagger}_{x}\sigma_{-}+g^{*}(x)a_{x}\sigma_{+}\right].$
@@
The equivalence between Eqs. (1) and (7) is guaranteed by the fact that the same effective action for spin degree of freedom can be derived from the Hamiltonians Eqs. (1) and (7) bulla1997 .
It is noteworthy that the forms of $h(x)$ or $g(x)$ are not uniquely determined for a given $J(x)$. By this freedom, one can choose proper form for $h(x)$ or $g(x)$ for the sake of discretization of the environment discretization . Given the Ohmic spectral function Eq. (2), one may choose
@@
The main idea of CDA is to utilize complex Gauss quadratures to discretize the continuum of the environment. For this purpose, $\int_{0}^{\infty}\text{d}x$ is replaced by a line integral $\int_{\Gamma}\text{d}z$ in the complex plane shown in Fig. 1, where $z=x+\mathbbm{i}y$. $\Gamma$ denotes the semicircle with radius $R$ in the lower half of the complex plane, centered at $\left(R,0\right)$. With the requirement that both $h(x)$ and $g(x)$ have no singularity in the complex plane, $\int_{\Gamma}\text{d}z$ is equal to $\int_{0}^{\infty}\text{d}x$ by Cauchy theorem if $R\rightarrow\infty$. Along the path $\Gamma$, $z=R(1+e^{\mathbbm{i}\theta})$ with $\theta\in\left[\pi,2\pi\right]$. Thus, the real part of $z$ is ensured to be nonnegative, characterizing the energy mode in the environment. The value of $R$ is responsible for the effective energy modes in the environment, used to simulate the spin dynamics.
Similar to the way in discretization , the following transformation is introduced:
@@
It is stressed that the inner product is defined deliberately without complex conjugation to construct the three-term recurrence relation for $\eta_{n}(z)$ gautschi . In this way, $\eta_{n}(z)$ can display two features, which are crucial for discretization of Eq. 7. One is the orthonormality defined as
@@
where $N$ is the highest degree of the polynomial, used in the evaluation. $h(z)$ and $g(z)$ can be obtained directly through replacing $x$ in Eq. (9) by $z$.
As a further simplification, let us introduce the new mode operator
@@
is presented in Fig. 2 for different values of $s$ and $\eta$ (the explicit plots for the evolution of survival probability are provided by Fig.A1 in Appendix I). Nonetheless, due to the time-consuming nature of solving Eq. (4), the exact evaluations are restricted up to $t=10$ with the step length of $10^{-4}$. For CDA, the weight function chosen is $w(z)=1$, which might cause the appearance of complex eigenmodes in the environment. It should be stressed that choosing $g(z)$ as the weight function will not generate a non-Hermitian effective Hamiltonian. Evidently, the computational errors has a magnitude of $10^{-3}$ in most circumstances, and thus CDA can provides reliable evaluation of the spin dynamics.
II.3 Introduction of $\tilde{H}_{\text{dis}}$
It is noted that Fig. 2(c) shows a significant increase of error when $s=1$. The explicit calculations reveals a slow convergence to a stable result as $N$ increases. To capture the long-term behavior of spin dynamics, we propose to use